OpenStudy (civicsiscool44):

The function f(t) = t2 + 12t − 18 represents a parabola. Part A: Rewrite the function in vertex form by completing the square. Show your work. (6 points) Part B: Determine the vertex and indicate whether it is a maximum or a minimum on the graph. How do you know? (2 points) Part C: Determine the axis of symmetry for f(t). (2 points)

OpenStudy (civicsiscool44):

@ganeshie8

OpenStudy (trojanpoem):

(t+6)^2 - 36-18 = y y = (t+6)^2 - 54 center is (-6, -54)  vertex is (h,k+a ) is ( -6, -54 + 1/4) axis of symmetry is y = -54. x = -6